Sol Summary – April 24th

Report Title: Sol Summary Report

Crew #: 335
Position: Crew Commander
Report Prepared by: David Laude
Date: 24 Apr 2026
Sol: 5

Summary Title: Kat’s Day Has Come

Mission Status: Steady and productive.

Sol Activity Summary: Today the crew’s rhythm was more casual than the usual two EVAs as we took Kat on her first easy EVA along with Commander and film makers, Elisa and Matteo after a morning of work on our various research projects on campus. Kat has had some medical issues and so we "tested" her on an easy, but beautiful drive most of the way to Brahe Highway since she was now feeling up to it. We stopped for Dave and Kat to practice piloting a drone while in EVA and the film makers were more than delighted to film us, drone and the intense Marian land scape. She had some pain about half way through when finishing the drone flights, but not enough in her opinion to return to base so she mostly rested in Opportunity. Tonight’s dinner is a much anticipated pizza.

Anomalies in Work: We still find headset #2 non-functional and so we put it aside.

Weather: Conditions were nearly ideal with 70F, light winds and some cloud cover.

Crew Physical Status: All crew members are generally in good condition. Some are experiencing mild symptoms such as sore throats and nasal congestion.

EVA: See above

GreenHab Report – April 24th

Report title: GreenHab Report

Crew #: 335

Position: GreenHab officer

Report prepared by: Charlotte Brugge

Date: 24APR2026

Sol: day 5

Environmental control (fan & heater): Automated

Average temperatures (last 24h): 91.9

Maximum temperature (last 24h): 114.8

Minimum temperature (last 24h): 69.0

Hours of supplemental light: Automated

Daily water usage for crops: 6 liters

Daily water usage for research and/or other purposes: 20 liters for the clay from yesterday night and today. The clay had to be rehydrated and filtered. Then to separate the clay from the silt, water is added to a bucket. You stir for 30/60 seconds and pour the water in another bucket. The clay particles will stay floating and the silt will rest at the bottom.

Water in the Tank (160 gal useful capacity): 102.4

Time(s) of watering for crops: 17.30

Narrative: watered and checked the plants, we filmed with the crew and cut off some flowers so the arugula can grow better

Harvest: (include which crop and mass in grams) 636 tomatoes (4) and basil

Support/supplies needed:0

Future GreenHab Improvement Observations: It would be good if all of the plants were properly labelled.

Journalist Report – April 24th

Crew #: 335

Position: Crew Journalist

Report Prepared by: Matteo Calore

Date: 24Apr2026

Sol: 5

Summary Title: Expansive Silence

This morning, most of us woke up later than usual. The night had been restless, and many crew members slept poorly. Our windowless sleeping cells make it difficult for some to breathe deeply, resulting in uneasy sleep and dense, vivid dreams.

In the morning, each of us took time to focus on our individual projects, while some crew members cleaned the Hab. Coordinating shared domestic tasks is not always easy—these are the ordinary negotiations of cohabitation. At times, some find themselves taking on more responsibilities than others.

In the afternoon, part of the crew went out for an EVA. Kat wanted to test the drone, and the film crew followed her to document the moment. The landscapes we encounter remain breathtaking: vast stretches of pink terrain in shifting tones, rocks shaped by time and wind, and silence extending for kilometers across these expanses. Here, the realm of rocks feels sovereign.

We returned to the Hab, and everyone resumed their respective tasks. In the evening, we prepared pizza. We harvested some cabbage from the greenhouse, which Charlotte has been consistently tending. The plants appear somewhat strained, but they remain alive and resilient.

Sol Summary – April 23rd

Report Title: Sol Summary Report

Crew #: 335
Position: Crew Artist
Report Prepared by: Elisa Strinna
Date: 23 Apr 2026
Sol: 4

Summary Title: Candor Chasma

Mission Status: Steady and productive despite environmental constraints.

Sol Activity Summary:Today the crew’s rhythm was slightly slower, as several members felt tired. We woke up between 6:00 and 7:00, had a short breakfast, and planned to hold our morning meeting at 8:00. However, the film crew encountered technical issues with the microphones, which required us to postpone the meeting until 12:00.

Once the issue was resolved, we held a brief meeting to review the day’s program and had lunch before preparing for the EVA. Due to earlier delays and additional organizational time, we exited approximately 30 minutes later than scheduled.

The EVA team—David, Charlotte, Elisa, and Matteo—headed to Candor Chasma, while Kat remained at the Hab to work on her astronomy project. During the EVA, Charlotte collected clay and David took photographs, while Elisa and Matteo documented Charlotte’s harvesting process.

We returned around 17:00. Afterward, each crew member took some time to rest and continue their individual projects. Dinner was held around 20:00, and we went to bed at approximately 22:00.

Look Ahead Plan: An EVA is planned for tomorrow afternoon, including Kat’s first EVA. Internal project work will continue in parallel.

Anomalies in Work:Some issues occurred with sound recording equipment used by the film crew.

Weather:Conditions were stable in the afternoon, with strong winds in the morning.

Crew Physical Status:All crew members are generally in good condition. Some are experiencing mild symptoms such as sore throats and nasal congestion.

EVA:One EVA was conducted to Candor Chasma. A second planned EVA did not take place due to strong winds in the morning and was postponed.

EVA Report – April 27th

EVA Report

Crew #: 335
Position: Elisa Strinna (Crew Artist)
EVA #: 11
Date: 27APR2026
Sol: 8

Purpose of EVA: The purpose of the EVA changed due to strong winds and the health condition of one crew member. Only two members of the film crew went outside to capture landscape footage around MDRS, including shots of the station at sunset.

Weather Conditions: Weather conditions supported the EVA with caution.
No precipitation was expected, with temperatures around 20°C (68°F). However, winds of 20–25 MPH impacted conditions.

Start Time: 17:50
End Time: 19:50

Narrative: We went outside to film landscapes around MDRS. We did not travel far due to the strong wind and the late timing. We captured images of MDRS from a distance along Cow Dung Road, then continued slightly further to film the surrounding area, which appears very Mars-like. After completing the filming, we returned to the station.

Destination: MDRS surroundings (UTM WGS84 not applicable)

Route: Cow Dung Road

Mode of Travel: Rover + Walking

EVA Participants: Elisa Strinna (Crew Artist), Matteo Calore (Crew HSO / Journalist)

Astronomy Report – April 27th

[category  astronomy-report]

Report title: Astronomy Report

Crew #: 335

Position: Engineer

Report prepared by: Katherine Berry

Date: 27Apr2026

Sol: 8

MDRS ROBOTIC OBSERVATORY

Robotic Telescope Requested: MDRS-WF

Objects to be Imaged this Evening: None, bad weather.

Images submitted with this report: None, images in process.

Problems Encountered: Mount error for the MDRS–WF

MUSK OBSERVATORY

Solar Features Observed: N/A

Images submitted with this report: N/A

Problems Encountered: N/A

Operations Report – April 27th

Report title: Operations Report

Crew #: 335

Position: Engineer

Report prepared by: Katherine Berry

Date: 27Apr2026

Sol: 8

Non-nominal systems: None

Notes on non-nominal systems: None

ROVERS

Spirit rover used:

Hours: 309.7

Beginning charge: 100%

Ending charge: 47%

Currently charging: Yes

Opportunity rover:

Hours: 0

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Curiosity rover used:

Hours: 0

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Perseverance rover used:

Hours: 0

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

General notes on rovers: None

Summary of Hab operations:

Water Use (see notes): 24.77gal/d

Main tank level (remaining gallons, see notes): 112 gal

Main water tank pipe heater (ON or OFF): OFF

Main water tank heater (ON or OFF) OFF (was removed)

Toilet tank emptied (NO or YES): NO

Summary of internet: Connection was used for research, accessing, and working on homework, sending reports, video editing and listening to music.

Summary of suits and radios: None

Summary of GreenHab operations: (use information from the GreenHab report)

Water use: 6 gallons

Heater (ON or OFF): Auto

Supplemental light (hours of operation): Auto

Harvest (name, weight in grams): 0g

Summary of ScienceDome operations: Used to film creative projects, drying clay and manipulating clay

Dual split (Heat or AC, On or Off or Automatic): Automatic

Summary of RAM operations: None

Summary of any observatory issues: Mount error for the MDRS–WF

Summary of health and safety issues: Crew member fell down stairs, monitoring.

Questions, concerns and requests to Mission Support: None

GreenHab Report – April 27th

Report title: GreenHab Report

Crew #: 335

Position: GreenHab officer

Report prepared by: Charlotte Brugge

Date: 27PR2026

Sol: day 7

Environmental control (fan & heater): Automated

Average temperatures (last 24h): 90.4

Maximum temperature (last 24h): 111.2

Minimum temperature (last 24h): 69.6

Hours of supplemental light: Automated

Daily water usage for crops: 6

Daily water usage for research and/or other purposes: 0

Water in the Tank (160 gal useful capacity): 89.6

Time(s) of watering for crops: 19.00

Narrative: watered and checked the plants.

Harvest: (include which crop and mass in grams) 0

Support/supplies needed:0

Future GreenHab Improvement Observations: None noticed today